Find two consecutive odd integers such that twice the larger is 17 more than the smaller.
Answer by mananth(16949)

let the odd integers be x & x+2
..
twice the larger = 2(x+2)
17 more than the smaller
= x+17
..
2(x+2) = x+17
2x+4 =x+17
2x-x= 17-4
x= 13
13, 15 are the integers
..
m.ananth@hotmail.ca